# In your own words: computationally identifying interpretable themes from free-text survey data

### 27 Mar 2026 v1

@@ Abstract
Free-text survey responses can provide nuance often missed by structured questions, but remain difficult to statistically analyze. To address this, we introduce In Your Own Words, a computational framework for exploratory analyses of free-text survey data that identifies structured, interpretable themes in free-text responses more precisely than previous computational approaches, facilitating systematic analysis. To illustrate the benefits of this approach, we apply it to a new dataset of free-text descriptions of race, gender, and sexual orientation from 1,004 U.S. participants. The themes our approach learns have three practical applications in survey research. First, the themes can suggest structured questions to add to future surveys by surfacing salient constructs—such as belonging and identity fluidity—that existing surveys do not capture. Second, the themes reveal heterogeneity within standardized categories, explaining additional variation in health, well-being, and identity importance. Third, the themes illuminate systematic discordance between self-identified and perceived identities, highlighting mechanisms of misrecognition that existing measures do not reflect. More broadly, our framework can be deployed in a wide range of survey settings to identify interpretable themes from free text, complementing existing qualitative methods.
@@ 1 Introduction
Addressing this, we introduce In Your Own Words, a computational framework for exploratory analyses of survey data that complements existing qualitative approaches. Our framework learns interpretable, human-readable themes from free-text responses by leveraging the powerful abilities of sparse autoencoders (SAEs) and large language models (LLMs), inspired by recent work showing that SAEs can successfully identify interpretable themes in text corpora [hypothesaes, movva2025s, jiangsun2025interp_embed]. Our approach does not require pre-specifying themes or manually annotating text; instead, it identifies recurring themes in how respondents describe themselves and annotates each free-text response with whether it contains each theme. Each response is then represented as a vector of the free-text themes it contains, preserving the richness and nuance of open-ended free text while converting the data into a structured form that can be easily used in downstream statistical analysis. Our work builds on a long tradition of classical and modern methods for computational text analysis [blei2003latent, grootendorst2022bertopic, pham-etal-2024-topicgpt, zhong2025hicode]; we show our approach outperforms several established methods from this literature.
@@ 2.2 Computational approach
Several conceptual features of this approach are worth highlighting. First, unlike computational methods that classify individuals into a single cluster or group, our approach allows each response to reflect multiple themes at once, acknowledging that identities are often multifaceted and overlapping [crenshaw2013demarginalizing]. Second, our approach does not require the themes to be pre-specified, allowing it to capture unexpected, context-specific nuances rather than relying on predefined hypotheses. Third, the approach is scalable: because it can be performed computationally and does not require hand-coding of the entire dataset, it can be applied even to datasets with millions of free-text responses. Finally, we emphasize that the goal of computationally learning themes from the data is to complement, not replace, qualitative coding methods, and specifically to facilitate the exploratory analyses we discuss in greater detail below.
@@
We also assess recall—confirming, in other words, that the computationally identified free-text themes do not miss salient motifs in the free-text responses. We perform a manual review of 100 randomly sampled responses per identity axis (race, gender, and sexual orientation). We find that, while our framework does miss some motifs, they generally occur rarely (1–10% of sampled responses). This is expected: because we constrain our method to identify only free-text themes, it prioritizes frequently-occurring motifs and therefore omits some low-frequency motifs. Manual validation suggests two main reasons motifs are missing. First, some are close variants of content already represented by our free-text themes, but expressed using adjacent labels or a different level of abstraction—for example, Hispanic/Latino identification, a missed motif, partially aligns with an existing free-text theme about speaking Spanish being part of cultural identity. Second, a smaller set of motifs reflect truly distinct content from those we capture with , including motifs discussing discrimination (5%) and racial privilege (5%) for race; feminism (1%) and fear or discomfort discussing gender (1%) in gender; and dating (9%) in sexual orientation. Importantly, however, we show that many of these motifs can be captured simply by increasing the number of free-text themes our framework learns (e.g., to ); we opt against doing this in our main analysis because it also produces more semantically overlapping or split themes, highlighting a tradeoff between coverage and redundancy. These findings underscore the value of complementing our computationally identified free-text themes with human-in-the-loop qualitative validation: manual review clarifies how the SAE and LLM represent and interpret responses and can pinpoint motifs that do not surface at an initial choice of . Different analytic use cases may motivate different choices of , or incorporating themes learned at multiple values of ; our framework can flexibly accommodate all these choices. We report detailed examples and analyses of these patterns in Appendix C.1.
@@ 2.3 Analysis of free-text themes
Figure 2 presents a selection of the free-text themes for race, gender, and sexual orientation (Figures S7–S9 plot the full set of themes). While some themes are expressed predominantly by members of a single standardized category—e.g., most respondents who express the “mentions being a male in a direct and explicit manner” theme are cisgender men—many themes cut across standardized categories rather than being expressed primarily by members of one category. The cross-cutting race themes, for example, capture cultural and social dimensions not represented by standardized categories, such as shared values and traditions (e.g., “describes beliefs or values associated with their race” or “mentions cultural food as important to identity”), language (“mentions speaking or understanding multiple languages”), immigration (“mentions being an immigrant or immigrant generation”), and belonging (“mentions a lack of racial or ethnic belonging”). For gender and sexual orientation, we similarly observe many cross-cutting themes: for example, themes mentioning questioning or struggling with gender identity or sexual orientation and themes mentioning fluidity. We also observe a noticeable divide between the themes expressed by non-minority222For gender, we define non-minority respondents as cisgender men and cisgender women; for sexual orientation, we define non-minority respondents as straight participants; for race, we define non-minority respondents as White participants. and minority respondents: for example, non-minority gender respondents were more likely to emphasize gender roles or responsibilities, while minority respondents were more likely to discuss fluidity in identity (both ).
@@ 3 Discussion
Looking ahead, the In Your Own Words framework offers substantial opportunities to inform the design, iteration, and analysis of surveys. The framework can be applied to analyze responses after a survey is completed, and can also support an iterative survey design process in which free-text responses are used to identify relevant features, guide the development of new structured questions, and evaluate whether existing measures adequately capture respondents’ lived experiences. This is particularly useful in settings where key concepts are evolving, theory is underdeveloped, or standardized measures lag behind cultural shifts in terminology or social norms. Importantly, we view the framework as a complement to qualitative inquiry rather than a replacement: it can surface interpretable patterns at scale that help prioritize where deeper reading, coding, or interviews are most needed, and can support mixed-method workflows for validating and refining constructs. While we demonstrate the benefits of our approach in the context of identity, the computational method could be applied much more generally to free-text survey questions in many domains (e.g., workplace climate, healthcare, or public opinion). Future work could extend this framework to incorporate free-text responses into theory-building and iterative survey design, enabling surveys that remain interpretable, scalable, and responsive to how people describe their experiences over time.
@@ Methods
Each participant answered three free-text questions asking them to describe their race, gender, and sexual orientation in their own words. The main prompt for each question read: “In at least 2–3 sentences, how would you describe your [identity]?” Each question was followed by a short italicized sub-prompt that provided examples of topics participants might discuss, such as cultural practices, life experiences, or identity-related milestones. We included these sub-prompts to encourage more detailed responses and mitigate the tendency toward brief, low-effort answers in online open-ended formats [aspinall2012answer], as prior work shows that such prompts can improve response length and quality [smyth2009open]. For race, the sub-prompt read: “For example, you could discuss specific traditions or customs, cultural practices or norms, languages or dialects spoken in your family, or ways race and/or ethnicity has influenced your life experiences.” The gender identity and sexual orientation questions followed the same format, with tailored examples suggesting that participants might describe milestones, social perceptions, or aspects of daily life related to that dimension of identity. The sub-prompts were grounded in established frameworks for each identity axis: the race prompt emphasized cultural norms, practices, language, and shared experiences, consistent with frameworks that define racial and ethnic identity through shared behaviors and values [phinney1996we], while the gender and sexual orientation prompts focused on narrative identity and development, aligning with life-story models of identity formation [mcadams2001psychology]. Table 1 provides full question wordings. As with any guided prompt, these sub-prompts may have shaped the range of responses we received—for instance, themes related to language, ancestry, and immigration emerged frequently across identity axes. Our computational framework is flexible; researchers who wish to focus on different identity-related themes could apply the same methodology using alternative prompts.
@@
In addition to the free-text identity questions, each participant selected from detailed multiple-choice options describing their race, gender, and sexual orientation. The demographic breakdown of participants is shown in Table S1. For race, each participant could select multiple categories from the following set: “American Indian or Alaska Native”, “Asian”, “Black or African American”, “Hispanic or Latino”, “Middle Eastern or North African”, “Native Hawaiian or Pacific Islander”, “White”, and “Some Other Race”. For gender, each participant first selected their current gender identity (“Man”, “Woman”, or “Some other way”), then also indicated whether they identify as transgender by selecting “Yes”, “No”, or “I prefer not to answer”. For sexual orientation, each participant could select one or more of the following: “Asexual or aromantic”, “Bisexual”, “Demisexual”, “Gay”, “Lesbian”, “Pansexual”, “Queer”, “Questioning”, “Sexually fluid”, “Straight or heterosexual”, “Other sexual identity or orientation (please specify)”, and “I prefer not to answer”. In Figures 2 and S7-S9, we visualize the concordance between standardized identity categories and free-text themes. To construct these visualizations, we create mutually exclusive identity categories from the original raw responses, which allow participants to select multiple categories; the details of this procedure are described in Appendix D.2. Aside from these figures and one supplementary analysis (Appendix B), all other analyses use the original multiple-choice data without regrouping them into mutually exclusive categories.
@@
Life satisfaction. To assess life satisfaction, we used the Cantril Self-Anchoring Ladder, a 0–10 scale in which participants indicate where they currently stand between the ‘worst possible life’ (0) and the ‘best possible life’ (10) [kilpatrick1960self]. This scale is a widely used measure of global well-being in large-scale population surveys (e.g., Gallup World Poll) [tortora2010gallup].
@@
The goal of this step is to annotate each free-text response for whether it contains each theme. We use GPT-4.1-mini for this task, following prior work demonstrating that LLMs can serve as reliable annotators in computational social science settings [asirvatham2026gpt]. The LLM is prompted separately for each theme–response pair and outputs a binary label indicating whether the theme is present (1) or absent (0) in the response. The annotation prompt is identical to the one used to measure fidelity in the LLM interpretation step (Figure S5). This procedure creates a binary matrix whose rows are free-text responses, whose columns are themes, and whose entries indicate whether a given response displays a given theme. This matrix is the final structured output of the computational framework and is used in the downstream quantitative analyses.
@@
We conduct linear regressions to assess whether free-text identity themes provide additional explanatory power for self-reported health and social outcomes beyond standardized identity categories. We consider six outcomes (described in §M.2): mental health, physical health, life satisfaction, everyday discrimination, income, and identity importance. Because these outcomes are measured on different scales (e.g., life satisfaction on a 0–10 scale and mental health on a 5-point ordinal scale), all outcome variables () are z-scored prior to analysis—that is, each outcome is transformed to have mean zero and standard deviation one within the analytic sample. This normalization is a standard transformation [angrist2009mostly] that allows regression coefficients to be interpreted in standard deviation units, making them comparable across outcomes. Categorical outcomes are first mapped to ordered numeric scales prior to standardization. All independent variables are binary indicators and are not rescaled.
@@
To better understand which specific free-text themes are associated with life outcomes, for each identity–outcome pair where adding free-text themes significantly improved model fit in the nested F-test, we estimate follow-up regression models that include each theme individually, alongside controls for the standardized identity categories for the corresponding identity axis. That is, for each free-text theme , we estimate the following model:
This approach helps clarify which individual themes drive the improvement in model fit observed in the nested F-tests. For example, we estimate a model where the outcome is self-reported mental health and the covariates include a binary indicator for whether the free-text response mentions “feelings of not belonging”, together with indicators for the respondent’s standardized race categories. We estimate models for each theme individually because themes may be collinear, complicating interpretation of multivariate models. We report point estimates and 95% confidence intervals (uncorrected for multiple hypotheses) for these models in Figures S10–S12. Due to the large number of themes, these analyses should be viewed as exploratory, as discussed in the main text.
@@ Appendix A Distribution of responses to the survey
Word count distribution: Most participants wrote relatively concise responses to the race, gender, and sexual orientation free-text responses, as shown in Figure S1. The median word count was 31 for race, 26 for gender, and 24 for sexual orientation. While a few individuals wrote longer responses (maximum: 282 for race, 272 for gender, and 352 for sexual orientation), the vast majority of responses were under 50 words, corresponding to the 75-85th percentile across all three axes. These distributions are similarly shaped—right-skewed with a long tail—and suggest that, despite the 2–3 sentence prompt, most participants did not write extensive narratives. These patterns, alongside our themes, underscore that our approach is operationalizable on responses of varying lengths. Figure S2 presents word counts for the perceived identity free-text responses, which tended to be shorter than the self-described identity responses. The median word count was 21 for race, 14 for gender, and 16 for sexual orientation, again suggesting that participants in our sample tended to have more to say about their race than their gender and sexual orientation.
@@
| Cisgender Man | 28.6% |
| Cisgender Other | 5.4% |
| Cisgender Woman | 49.5% |
@@
| Other sexual identity or orientation please specify | 2.2% |
@@
In Figure S3 (right), we compare the relative length of free-text identity responses across categories. Again, we observe meaningful heterogeneity within each identity axis. Among race categories, American Indian or Alaska Native, Hispanic or Latino, Middle Eastern or North African, and multiracial respondents wrote significantly more than White respondents. Asian and Native Hawaiian or Pacific Islander individuals did not differ significantly, and Black or African American respondents wrote slightly fewer words on average (a small but statistically significant difference). Within the gender responses, minorities besides transgender women wrote significantly more than cisgender men and women. For sexual orientation, minority categories—including asexual, bisexual, gay or lesbian, and queer respondents—wrote more than heterosexual respondents. These findings show that respondents with minority identities were often among those who wrote the most and valued the opportunity to describe themselves in their own words.

### 03 Apr 2026 v2

@@ Abstract
Free-text survey responses can provide nuance often missed by structured questions, but remain difficult to statistically analyze. To address this, we introduce In Your Own Words, a computational framework for exploratory analyses of free-text survey data that identifies structured, interpretable themes in free-text responses, facilitating systematic analysis. To illustrate the benefits of this approach, we apply it to a new dataset of free-text descriptions of race, gender, and sexual orientation from 1,004 U.S. participants. The themes our approach produces on this dataset are more coherent and interpretable than those produced by past computational methods. The themes have three practical applications in survey research. First, they can suggest structured questions to add to future surveys by surfacing salient constructs—such as belonging and identity fluidity—that existing surveys do not capture. Second, the themes reveal heterogeneity within standardized categories, explaining additional variation in health, well-being, and identity importance. Third, the themes illuminate systematic discordance between self-identified and perceived identities, highlighting mechanisms of misrecognition that existing measures do not reflect. More broadly, our framework can be deployed in a wide range of survey settings to identify interpretable themes from free text, complementing existing qualitative methods.
@@ 1 Introduction
Addressing this, we introduce In Your Own Words, a computational framework for exploratory analyses of survey data that complements existing qualitative approaches. Our framework learns interpretable, human-readable themes from free-text responses by leveraging the powerful abilities of sparse autoencoders (SAEs) and large language models (LLMs), inspired by recent work showing that SAEs can successfully identify interpretable themes in text corpora [hypothesaes, movva2025s, jiangsun2025interp_embed]. Our approach does not require pre-specifying themes or manually annotating text; instead, it identifies recurring themes in how respondents describe themselves and annotates each free-text response with whether it contains each theme. Each response is then represented as a vector of the free-text themes it contains, preserving the richness and nuance of open-ended free text while converting the data into a structured form that can be easily used in downstream statistical analysis. Our work builds on a long tradition of classical and modern methods for computational text analysis [blei2003latent, grootendorst2022bertopic, pham-etal-2024-topicgpt, zhong2025hicode]; we show that our approach outperforms several methods from this literature on the data we collect.
@@ 2.2 Computational approach
Several conceptual features of this approach are worth highlighting. First, unlike computational methods that classify individuals into a single cluster or group, our approach allows each response to reflect multiple themes at once, acknowledging that identities are often multifaceted and overlapping [crenshaw2013demarginalizing]. Second, our approach does not require the themes to be pre-specified, allowing it to capture unexpected, context-specific nuances rather than relying on predefined hypotheses. Third, the approach is scalable: because it can be performed computationally and does not require hand-coding of the entire dataset, it can be applied even to datasets with millions of free-text responses. Finally, we emphasize that the goal of computationally learning themes from the data is to complement, not replace, qualitative coding methods [nelson2020computational], and specifically to facilitate the exploratory analyses we discuss in greater detail below.
@@
We also assess recall—confirming, in other words, that the computationally identified free-text themes do not miss salient motifs in the free-text responses. We perform a manual review of 100 randomly sampled responses per identity axis (race, gender, and sexual orientation). We find that, while our framework does miss some motifs, they generally occur rarely (1-10% of sampled responses). This is expected: because we constrain our method to identify only free-text themes, it prioritizes frequently-occurring motifs and therefore omits some low-frequency motifs. Manual validation suggests two main reasons motifs are missing. First, some are close variants of content already represented by our free-text themes, but expressed using adjacent labels or a different level of abstraction—for example, Hispanic/Latino identification, a missed motif, partially aligns with an existing free-text theme about speaking Spanish being part of cultural identity. Second, a smaller set of motifs reflect truly distinct content from those we capture with , including motifs discussing discrimination (5%) and racial privilege (5%) for race; feminism (1%) and fear or discomfort discussing gender (1%) in gender; and dating (9%) in sexual orientation. Importantly, however, we show that many of these motifs can be captured simply by increasing the number of free-text themes our framework learns (e.g., to ); we opt against doing this in our main analysis because it also produces more semantically overlapping or split themes, highlighting a tradeoff between coverage and redundancy. These findings underscore the value of complementing our computationally identified free-text themes with human-in-the-loop qualitative validation: manual review clarifies how the SAE and LLM represent and interpret responses and can pinpoint motifs that do not surface at an initial choice of . Different analytic use cases may motivate different choices of , or incorporating themes learned at multiple values of ; our framework can flexibly accommodate all these choices. We report detailed examples and analyses of these patterns in Appendix C.1.
@@ 2.3 Analysis of free-text themes
Figure 2 presents a selection of the free-text themes for race, gender, and sexual orientation (Figures S7-S9 plot the full set of themes). While some themes are expressed predominantly by members of a single standardized category—e.g., most respondents who express the “mentions being a male in a direct and explicit manner” theme are cisgender men—many themes cut across standardized categories rather than being expressed primarily by members of one category. The cross-cutting race themes, for example, capture cultural and social dimensions not represented by standardized categories, such as shared values and traditions (e.g., “describes beliefs or values associated with their race” or “mentions cultural food as important to identity”), language (“mentions speaking or understanding multiple languages”), immigration (“mentions being an immigrant or immigrant generation”), and belonging (“mentions a lack of racial or ethnic belonging”). For gender and sexual orientation, we similarly observe many cross-cutting themes: for example, themes mentioning questioning or struggling with gender identity or sexual orientation and themes mentioning fluidity. We also observe a noticeable divide between the themes expressed by non-minority222For gender, we define non-minority respondents as cisgender men and cisgender women; for sexual orientation, we define non-minority respondents as straight participants; for race, we define non-minority respondents as White participants. and minority respondents: for example, non-minority gender respondents were more likely to emphasize gender roles or responsibilities, while minority respondents were more likely to discuss fluidity in identity (both ).
@@ 3 Discussion
Looking ahead, the In Your Own Words framework offers substantial opportunities to inform the design, iteration, and analysis of surveys. The framework can be applied to analyze responses after a survey is completed, and can also support an iterative survey design process in which free-text responses are used to identify relevant features, guide the development of new structured questions, and evaluate whether existing measures adequately capture respondents’ lived experiences. This is particularly useful in settings where key concepts are evolving, theory is underdeveloped, or standardized measures lag behind cultural shifts in terminology or social norms. Importantly, we view the framework as a complement to qualitative inquiry rather than a replacement: it can surface interpretable patterns at scale that help prioritize where deeper reading, coding, or interviews are most needed, and can support mixed-method workflows for validating and refining constructs [nelson2020computational]. While we demonstrate the benefits of our approach in the context of identity, the computational method could be applied much more generally to free-text survey questions in many domains (e.g., workplace climate, healthcare, or public opinion). Future work could extend this framework to incorporate free-text responses into theory-building and iterative survey design, enabling surveys that remain interpretable, scalable, and responsive to how people describe their experiences over time.
@@ Methods
Each participant answered three free-text questions asking them to describe their race, gender, and sexual orientation in their own words. The main prompt for each question read: “In at least 2-3 sentences, how would you describe your [identity]?” Each question was followed by a short italicized sub-prompt that provided examples of topics participants might discuss, such as cultural practices, life experiences, or identity-related milestones. We included these sub-prompts to encourage more detailed responses and mitigate the tendency toward brief, low-effort answers in online open-ended formats [aspinall2012answer], as prior work shows that such prompts can improve response length and quality [smyth2009open]. For race, the sub-prompt read: “For example, you could discuss specific traditions or customs, cultural practices or norms, languages or dialects spoken in your family, or ways race and/or ethnicity has influenced your life experiences.” The gender identity and sexual orientation questions followed the same format, with tailored examples suggesting that participants might describe milestones, social perceptions, or aspects of daily life related to that dimension of identity. The sub-prompts were grounded in established frameworks for each identity axis: the race prompt emphasized cultural norms, practices, language, and shared experiences, consistent with frameworks that define racial and ethnic identity through shared behaviors and values [phinney1996we], while the gender and sexual orientation prompts focused on narrative identity and development, aligning with life-story models of identity formation [mcadams2001psychology]. Table 1 provides full question wordings. As with any guided prompt, these sub-prompts may have shaped the range of responses we received—for instance, themes related to language, ancestry, and immigration emerged frequently across identity axes. Our computational framework is flexible; researchers who wish to focus on different identity-related themes could apply the same methodology using alternative prompts.
@@
In addition to the free-text identity questions, each participant selected from detailed multiple-choice options describing their race, gender, and sexual orientation. The demographic breakdown of participants is shown in Table S1. For race, each participant could select multiple categories from the following set: “American Indian or Alaska Native”, “Asian”, “Black or African American”, “Hispanic or Latino”, “Middle Eastern or North African”, “Native Hawaiian or Pacific Islander”, “White”, and “Some Other Race”. For gender, each participant first selected their current gender identity (“Man”, “Woman”, or “Some other way”), then also indicated whether they identify as transgender by selecting “Yes”, “No”, or “Prefer not to answer”. For sexual orientation, each participant could select one or more of the following: “Asexual or aromantic”, “Bisexual”, “Demisexual”, “Gay”, “Lesbian”, “Pansexual”, “Queer”, “Questioning”, “Sexually fluid”, “Straight or heterosexual”, “Other sexual identity or orientation (please specify)”, and “Prefer not to answer”. In Figures 2 and S7-S9, we visualize the concordance between standardized identity categories and free-text themes. To construct these visualizations, we create mutually exclusive identity categories from the original raw responses, which allow participants to select multiple categories; the details of this procedure are described in Appendix D.2. Aside from these figures and one supplementary analysis (Appendix B), all other analyses use the original multiple-choice data without regrouping them into mutually exclusive categories.
@@
Life satisfaction. To assess life satisfaction, we used the Cantril Self-Anchoring Ladder, a 0-10 scale in which participants indicate where they currently stand between the ‘worst possible life’ (0) and the ‘best possible life’ (10) [kilpatrick1960self]. This scale is a widely used measure of global well-being in large-scale population surveys (e.g., Gallup World Poll) [tortora2010gallup].
@@
The goal of this step is to annotate each free-text response for whether it contains each theme. We use GPT-4.1-mini for this task, following prior work demonstrating that LLMs can serve as reliable annotators in qualitative coding and related computational social science settings [than2025updating, asirvatham2026gpt]. The LLM is prompted separately for each theme-response pair and outputs a binary label indicating whether the theme is present (1) or absent (0) in the response. The annotation prompt is identical to the one used to measure fidelity in the LLM interpretation step (Figure S5). This procedure creates a binary matrix whose rows are free-text responses, whose columns are themes, and whose entries indicate whether a given response displays a given theme. This matrix is the final structured output of the computational framework and is used in the downstream quantitative analyses.
@@
We conduct linear regressions to assess whether free-text identity themes provide additional explanatory power for self-reported health and social outcomes beyond standardized identity categories. We consider six outcomes (described in §M.2): mental health, physical health, life satisfaction, everyday discrimination, income, and identity importance. Because these outcomes are measured on different scales (e.g., life satisfaction on a 0-10 scale and mental health on a 5-point ordinal scale), all outcome variables () are z-scored prior to analysis—that is, each outcome is transformed to have mean zero and standard deviation one within the analytic sample. This normalization is a standard transformation [angrist2009mostly] that allows regression coefficients to be interpreted in standard deviation units, making them comparable across outcomes. Categorical outcomes are first mapped to ordered numeric scales prior to standardization. All independent variables are binary indicators and are not rescaled.
@@
To better understand which specific free-text themes are associated with life outcomes, for each identity-outcome pair where adding free-text themes significantly improved model fit in the nested F-test, we estimate follow-up regression models that include each theme individually, alongside controls for the standardized identity categories for the corresponding identity axis. That is, for each free-text theme , we estimate the following model:
This approach helps clarify which individual themes drive the improvement in model fit observed in the nested F-tests. For example, we estimate a model where the outcome is self-reported mental health and the covariates include a binary indicator for whether the free-text response mentions “feelings of not belonging”, together with indicators for the respondent’s standardized race categories. We estimate models for each theme individually because themes may be collinear, complicating interpretation of multivariate models. We report point estimates and 95% confidence intervals (uncorrected for multiple hypotheses) for these models in Figures S10-S12. Due to the large number of themes, these analyses should be viewed as exploratory, as discussed in the main text.
@@ Appendix A Distribution of responses to the survey
Word count distribution: Most participants wrote relatively concise responses to the race, gender, and sexual orientation free-text responses, as shown in Figure S1. The median word count was 31 for race, 26 for gender, and 24 for sexual orientation. While a few individuals wrote longer responses (maximum: 282 for race, 272 for gender, and 352 for sexual orientation), the vast majority of responses were under 50 words, corresponding to the 75-85th percentile across all three axes. These distributions are similarly shaped—right-skewed with a long tail—and suggest that, despite the 2-3 sentence prompt, most participants did not write extensive narratives. These patterns, alongside our themes, underscore that our approach is operationalizable on responses of varying lengths. Figure S2 presents word counts for the perceived identity free-text responses, which tended to be shorter than the self-described identity responses. The median word count was 21 for race, 14 for gender, and 16 for sexual orientation, again suggesting that participants in our sample tended to have more to say about their race than their gender and sexual orientation.
@@
| Cisgender, Man | 28.6% |
| Cisgender, Woman | 49.5% |
| Cisgender, Another gender | 5.4% |
| Transgender, Man | 2.1% |
| Transgender, Woman | 1.8% |
| Transgender, Another gender | 9.8% |
@@
| Other sexual identity or orientation | 2.2% |
@@
In Figure S3 (right), we compare the relative length of free-text identity responses across categories. Again, we observe meaningful heterogeneity within each identity axis. Among race categories, American Indian or Alaska Native, Hispanic or Latino, Middle Eastern or North African, and multiracial respondents wrote significantly more than White respondents. Asian and Native Hawaiian or Pacific Islander individuals did not differ significantly, and Black or African American respondents wrote slightly fewer words on average (a small but statistically significant difference). Within the gender responses, minorities besides transgender women wrote significantly more than cisgender men and women. For sexual orientation, minority categories—including asexual, bisexual, gay or lesbian, and queer respondents—wrote more than heterosexual respondents. These findings suggest that respondents with minority identities were often among those who wrote the most and valued the opportunity to describe themselves in their own words.